# Black Mirrors: CPT-Symmetric Alternatives to Black Holes

### v1

@@ Abstract
Einstein’s equations imply that a gravitationally collapsed object forms an event horizon. But what lies on the other side of this horizon? In this paper, we question the reality of the conventional solution (the black hole), and point out another, topologically distinct solution: the black mirror. In the black hole solution, the horizon connects the exterior metric to an interior metric which contains a curvature singularity. In the black mirror, the horizon instead connects the exterior metric to its own CPT mirror image, yielding a solution with no curvature singularities. We give the general stationary (charged, rotating) black mirror solution explicitly, and also describe the general black mirror formed by gravitational collapse. The black mirror is the relevant stationary point when the quantum path integral is equipped with suitably CPT-symmetric boundary conditions, that we propose. It appears to avoid many vexing puzzles which plague the conventional black hole.
@@ I Introduction
The Einstein-Rosen bridge (the first wormhole geometry) was put forth by Einstein and Rosen in Einstein and Rosen (1935). One of their motivations was to get rid of the apparent “singularity” at the event horizon $r=2m$ by introducing a new coordinate $u$ by $r=2m+u^{2}$, and letting $u$ extend over $-\infty<u<\infty$. This describes two Schwarzschild exteriors (regions $I$ and $I^{\prime}$ in Fig. 1a), glued at the radius $r=2m$ (the midpoint in Fig. 1a). However, it was later deemed to be unphysical as the papers of Oppenheimer-Snyder Oppenheimer and Snyder (1939) describing spherical collapse showed that the singularity at $r=2m$ is just a coordinate singularity, and an infalling observer following the collapsing matter will not feel anything unusual happening at the event horizon. This (and other developments) led to the modern belief that the black hole interior really exists, and that the black hole’s horizon is completely non-singular.
@@
In this paper, we argue that certain natural modifications of Einstein-Rosen bridges – which we call black mirrors – are, in fact, direct consequences of CPT symmetry, and offer novel solutions to the problems mentioned above, along with other merits discussed below. [^1] We also draw the reader’s attention to several other proposals for variations to the usual black hole geometry or topology Gibbons (1986); Saravani et al. (2015); Visser (1989a, b); ’t Hooft (2016, 2017).
@@ II The Schwarzschild black mirror
For simplicity, let us first consider the Schwarzschild metric, describing an eternal, static, uncharged object. In Schwarzschild coordinates $(r,t)$ it is
@@
where
$f(r)\equiv 1-\frac{2m}{r}\quad{\rm and}\quad d\Omega^{2}\equiv d\theta^{2}+{\rm sin}^{2}\theta\,d\varphi^{2}.$ (2)
This solution has a horizon at $r=2m$, the root of $f(r)$. Note that Schwarzschild coordinates only cover the exterior region $r>2m$, and become singular at the horizon $r=2m$ (which lies at $t=\pm\infty$ in these coordinates).
The conventional view is that the full spacetime (including the horizon, and what lies behind it) is the maximal analytic extension obtained by switching from Schwarzschild to Kruskal coordinates d’Invemo (1992). The Penrose diagram of this extension is shown in Fig. 1a: it has two exterior regions $I$, $I^{\prime}$, two interior regions $II$, $II^{\prime}$, two black horizons (BHs) that particles can only enter, and two white horizons (WHs) that particles can only exit.
The simplest coordinate system that non-singularly covers both the exterior $I$ and its BH (WH) is the ingoing (outgoing) Eddington-Finkelstein system, in which we replace Schwarzschild time $t$ by ingoing (outgoing) Eddington-Finkelstein time $v_{+}$ ($v_{-}$), related by $dv_{\pm}=dt\pm\frac{dr}{f(r)}$, so that the line element (1) becomes
$ds^{2}=-f(r)dv_{\pm}^{2}\pm 2dv_{\pm}dr+r^{2}d\Omega^{2}.$ (3)
Figure 1: (a) Penrose diagram for the eternal Schwarzschild (Kruskal) black hole, indicating the two exteriors ($I$, $I^{\prime}$), two interiors ($II$, $II^{\prime}$), two black horizons (BH), two white horizons (WH), and the time-like killing vector $\partial_{t}$: time flows upward in exterior $I$, and downward in exterior $I^{\prime}$. (b) Penrose diagram for eternal Schwarzschild black mirror: exteriors I, I’ are glued together, BH to BH and WH to WH, as indicated (e.g. $P\sim P^{\prime}$, $R\sim R^{\prime}$), and with antipodal inversion of the $(\theta,\varphi)$ 2-sphere (see Appendix C), so the interiors II, II’ and their curvature singularities (purple regions) are removed.
@@
To answer this question, let us first take a step back to the conical geometry of Fig. 2a. Here it is clear that the single (blue) cone naturally extends to the double (blue+red cone) by letting $\sigma$ extend to negative values. Then, when we identify $\tau\sim\tau+8\pi m$ (removing the conical singularity), we obtain the full (blue+red) Euclidean Schwarzschild metric, Fig. 2b. Note: this full geometry still has nothing corresponding to the black hole interior; instead, it has two exteriors glued together at $\sigma=0$ (an $S^{2}$ of area $16\pi m^{2}$, corresponding to the horizon).
Further evidence for the two-sided geometry in Fig. 2b comes from black hole entropy. Gibbons and Hawking Gibbons and Hawking (1977) compute the entropy of the Schwarzschild black hole by computing the action of the Euclidean Schwarzschild metric (which comes entirely from the Gibbons-Hawking-York boundary term York (1972); Gibbons and Hawking (1977)). And, because they compute this action for the one-sided (blue) spacetime, they obtain a non-zero answer. This is a clue that the one-sided Euclidean metric is physically incomplete. If, instead, we compute the action/entropy for the two-sided Euclidean spacetime, the answer vanishes (as the two boundary terms, on isometric but oppositely oriented boundaries, make equal and opposite contributions, see Appendix B). This suggests a picture in which the full (blue+red) geometry corresponds to the full (pure, zero entropy) quantum state, while the half (blue) geometry corresponds to the mixed state obtained by tracing out the other (red) half; and the black hole entropy corresponds to the entanglement entropy (i.e. the von Neumann entropy of the blue side, after tracing out the red side). This in turn explains why the entropy is proportional to the area of the horizon (since the entanglement entropy between two regions is characteristically proportional to the size of the boundary separating them).
If we Wick rotate the full Euclidean metric (5) back to Lorentzian signature using $\tau=it$, we obtain a metric that now covers the full exterior $r>2m$ in Fig. 1a (i.e. regions $I$ and $I^{\prime}$). This is the Einstein-Rosen bridge Einstein and Rosen (1935), which is incomplete as it stands, and needs to be extended to a complete spacetime. The conventional completion is the Kruskal extension of Fig. 1a; but here we suggest a different completion. The simplest way to introduce this alternative possibility is to extend the metric to the horizon by replacing $t$, as before, with the ingoing (outgoing) Eddington-Finkelstein time coordinate $v_{+}$ ($v_{-}$). Then the metric becomes
@@
Now notice that, in order for this extension to be well-defined (to not assign the same $v,\sigma$ coordinate to two distinct points), we are forced (as shown in Fig. 1b) to identify the two BH horizons (e.g. $P\sim P^{\prime},Q\sim Q^{\prime}$) and similarly the two WH horizons (e.g. $R\sim R^{\prime},S\sim S^{\prime}$). [^2] Note that we do not identify all antipodal pairs in Fig. 1a, as has sometimes been considered in the past Gibbons (1986); in particular, we do not identify points in exterior $I$ with points in exterior $I^{\prime}$. This has the effect of gluing the exterior $I$ directly to the exterior $I^{\prime}$, which together form a complete spacetime.
The resulting spacetime is the simplest (Schwarzschild) black mirror: eternal, static, uncharged. It is described by the metric (7), with $-\infty<v<\infty$ and $-\infty<\sigma<\infty$. More precisely, the ingoing (outgoing) Eddington-Finkelstein coordinates $v_{+},\sigma$ ($v_{-},\sigma$) cover the whole spacetime except the WH (BH), and these two coordinate systems together form a complete atlas. We see that, in $v,\sigma$ coordinates, all components of the metric are explicitly smooth, analytic and finite functions of $v,\sigma$ (for all $-\infty<v,\sigma<\infty$); and one can check that all associated curvature tensors and invariants (such as the Kretschmann scalar) are as well.
Figure 3: The Penrose diagram for: (a) a black hole formed by gravitational collapse; or (b,c) a black mirror formed by gravitational collapse, in unfolded view (b) or folded view (c). Some infalling particle trajectories are also shown, with the angular $S^{2}$ over each point on the right (blue) side the antipodal inverse of the one on the left, see Appendix C.
Note that the Lorentzian and Euclidean geometries are now in perfect correspondence: in both, the $\sigma=0$ surface (i.e. the origin of the double cone in Euclidean signature, or the event horizon in Lorentzian signature) serves as the boundary between two mirror-image exterior regions, with no interior regions or curvature singularities.
But there is no free lunch! Although we have gotten rid of all curvature singularities, they have been replaced by another (milder) singularity. In particular, if we compute the eigenvalues of $g_{\mu\nu}$ (in $v,\sigma$ coordinates) then, in addition to the angular eigenvalues $\lambda_{\theta}=r^{2}(\sigma)$ and $\lambda_{\varphi}=r^{2}(\sigma){\rm sin}^{2}\theta$, we see the remaining two (positive and negative) eigenvalues
@@
Figure 4: (a) Traditional Penrose diagram for a black hole which collapses, then evaporates; (b) Improved Penrose diagram indicating shrinking of horizon due to Hawking radiation; (c) Black mirror which collapses, then evaporates.
@@ III Discussion
Figure 5: Asymptotically flat spacetime with CPT-symmetric boundary conditions: states $|\bar{i}\rangle$ and $|\bar{f}\rangle$ are the CPT conjugates of states $|i\rangle$ and $f\rangle$, respectively (see Appendix C).
So far, we have considered the general stationary black mirror. Now let us explain how we think the general time-dependent case works. (Since this part of the story can no longer be described explicitly analytically, it is correspondingly less certain and more speculative.) We start with the Penrose diagram in Fig. 3a for an asymptotically-flat black hole formed by realistic gravitational collapse, with a blue curve showing the trajectory of an infalling particle (the arrow shows the direction of flow of electric charge). The corresponding black mirror is again obtained by identifying two CPT mirror copies of the exterior at their respective horizons, as in Fig. 3b: now we see that the (blue) particle falling into the BH from one side meets its CPT mirror (red) anti-particle falling in from the other side. For visualization, it is convenient to fold the diagram in Fig. 3b to obtain the diagram shown in Fig. 3c. In this folded picture, time runs in its usual upward direction on both sheets, and we can explicitly see that both sheets share a common future null boundary (i.e., a BH).
In the stationary case, the event horizon and apparent horizon coincide. More generally, where they do not, we conjecture the correct matching surface is the apparent horizon (which is well-defined to linear order in perturbations around the stationary solution Hollands et al. (2024)) since: (i) it is defined locally (unlike the event horizon); and (ii) it appears to be the surface associated to dynamical black hole entropy Hollands et al. (2024) which, as argued above, is also naturally associated to the surface separating the two exteriors of a black mirror. This suggests the potential, at least in principle, of observationally distinguishing dynamical black mirrors from dynamical black holes, e.g. via gravitational wave observations.
@@
i) It is natural to consider the path integral for our universe with CPT-symmetric boundary conditions as shown in Fig. 6 (see Appendices B, C). Whereas the black hole (Fig. 1a) and black mirror (Fig. 1b) both satisfy these boundary condition, the black mirror seems to be the relevant saddle point as it (i) is free of curvature singularities, and (ii) has no additional boundaries (whereas the black hole solution has additional boundaries which are singular and also require further data to be specified).
@@
iii) The black hole information paradox Hawking (1976) and firewall paradox Almheiri et al. (2013) come from imagining that quantum information falls into a black hole, and asking how it can re-emerge (particularly since the black hole interior is causally disconnected from future null infinity, see Figs. 3a, 4a). In the black mirror Fig. 3c, with no interior, the puzzle is resolved: when a (blue) particle and (red) anti-particle meet at the horizon, they annihilate into (green) radiation that propagates along the horizon (classically), or gradually leaks off the horizon (quantumly), until the black hole evaporates (Fig. 4b).
Figure 6: The Penrose diagram for: (a) an unperturbed CPT-symmetric universe; a CPT-symmetric universe with two black mirrors.
iv) As a black hole settles to stationarity, we expect it to approach a thermal state, which should also be CPT-symmetric Sewell (1982). The black hole (Fig. 3a) does not do this (it settles down to a spacetime with a BH and no WH, with no way for CPT to act), but the black mirror (Fig. 3b, 4b) does (see Appendix C). (Hawking also argued that CPT conflicts with black hole interiors Hawking (2014).)
v) In the standard picture, one can drop global net charge into a black hole, which is then erased when the hole evaporates, leading to the lore that quantum gravity violates global symmetries Hawking (1975); Witten (2018); Harlow and Ooguri (2019, 2021). By contrast, in the black mirror, global charge falling in on one side is canceled by anti-charge falling in on the other; so black mirror evaporation does not imply violation of global symmetry. This is compatible with the fact that in the standard model $B-L$ is anomaly free and might be exact.
vi) This CPT-symmetric picture for black mirrors fits naturally in the CPT-symmetric picture of the cosmos advocated in Boyle et al. (2018, 2022a); Boyle and Turok (2021a, b); Boyle et al. (2022b); Turok and Boyle (2024); Boyle and Turok (2024); Turok and Boyle (2023) (see Fig. 6): the two sheets of spacetime split at the Bang, and merge at BHs. As before, it is natural to regard the two-sheeted spacetime as representing a pure state, while one sheet represents the mixed state obtained by tracing over the other sheet. This neatly explains why the corresponding entropy, associated with the boundary surface separating the two sheets, is given by the horizon area (in the BH case) Bekenstein (1972, 1973); Gibbons and Hawking (1977), and the total comoving spatial volume (in the cosmological case Turok and Boyle (2024); Boyle and Turok (2024)).

### v2

@@ Abstract
Einstein’s equations imply that a gravitationally collapsed object forms an event horizon. But what lies on the other side of this horizon? In this paper, we question the reality of the conventional solution (the black hole), and point out another, topologically distinct solution: the black mirror. In the black hole solution, the horizon connects the exterior metric to an interior metric which contains a curvature singularity. In the black mirror, the horizon instead connects the exterior metric to its own CPT mirror image, yielding a solution with smooth, bounded curvature. We give the general stationary (charged, rotating) black mirror solution explicitly, and also describe the general black mirror formed by gravitational collapse. The black mirror is the relevant stationary point when the quantum path integral is equipped with suitably CPT-symmetric boundary conditions, that we propose. It appears to avoid many vexing puzzles which plague the conventional black hole.
@@ I Introduction
The Einstein-Rosen bridge (the first wormhole geometry) was put forth by Einstein and Rosen in Einstein and Rosen (1935). One of their motivations was to get rid of the apparent “singularity” at the event horizon $r=2m$ by introducing a new coordinate $u$ by $r=2m+u^{2}$, and letting $u$ extend over $-\infty<u<\infty$. This describes two Schwarzschild exteriors (regions $I$ and $I^{\prime}$ in Fig. 1), glued at the radius $r=2m$ (the midpoint in Fig. 1). However, it was later deemed to be unphysical as the papers of Oppenheimer-Snyder Oppenheimer and Snyder (1939) describing spherical collapse showed that the singularity at $r=2m$ is just a coordinate singularity, and an infalling observer following the collapsing matter will not feel anything unusual happening at the event horizon. This (and other developments) led to the modern belief that the black hole interior really exists, and that the black hole’s horizon is completely non-singular.
@@
In this paper, we argue that certain natural modifications of Einstein-Rosen bridges – which we call black mirrors – are, in fact, direct consequences of CPT symmetry, and offer novel solutions to the problems mentioned above, along with other merits discussed below. [^1] Also note several other proposals for variations to the usual black hole geometry or topology Gibbons (1986); Saravani et al. (2015); Visser (1989a, b); ’t Hooft (2016, 2017); Carroll et al. (2009); Loran and Sheikh-Jabbari (2010, 2011); Koiran et al. (2024); Bouhmadi-López et al. (2020); Brahma et al. (2022) (and see Appendix B).
@@ II The Schwarzschild Black Mirror
For simplicity, let us first consider the Schwarzschild metric, describing an eternal, static, uncharged object. In Schwarzschild coordinates $(t,r,\theta,\varphi)$ it is
@@
where $f(r)\equiv 1-(2m/r)$ and $d\Omega^{2}\equiv d\theta^{2}+{\rm sin}^{2}\theta\,d\varphi^{2}$. This solution has a horizon at $r=2m$, the root of $f(r)$. Note that Schwarzschild coordinates only cover the exterior region $r>2m$, and become singular at the horizon $r=2m$ (which lies at $t=\pm\infty$ in these coordinates).
The conventional view is that the full spacetime (including the horizon, and what lies behind it) is the maximal analytic extension obtained by switching from Schwarzschild to Kruskal coordinates d’Invemo (1992). The Penrose diagram of this extension is shown in Fig. 1: it has two exterior regions $I$, $I^{\prime}$, two interior regions $II$, $II^{\prime}$, two black horizons (BHs) that particles can only enter, and two white horizons (WHs) that particles can only exit.
Figure 1: Penrose diagram for the eternal Schwarzschild (Kruskal) black hole, indicating the two exteriors ($I$, $I^{\prime}$), two interiors ($II$, $II^{\prime}$), two black horizons (BH), two white horizons (WH), and the time-like killing vector $\partial_{t}$, which points upward in exterior $I$, and downward in exterior $I^{\prime}$.
@@
To answer this question, let us first take a step back to the conical geometry of Fig. 2a. Here it is clear that the single (blue) cone naturally extends to the double (blue+red cone) by letting $\sigma$ extend to negative values, just as Einstein and Rosen did Einstein and Rosen (1935) on the Lorentzian side, with their coordinate $u=\sigma/\sqrt{8m}$. Then, when we identify $\tau\sim\tau+8\pi m$ (removing the conical singularity), we obtain the full (blue+red) Euclidean Schwarzschild metric, Fig. 2b. Note: this full geometry still has nothing corresponding to the black hole interior; instead, it has two exteriors glued together at $\sigma=0$ (an $S^{2}$ of area $16\pi m^{2}$, corresponding to the horizon).
Further evidence for the two-sided geometry in Fig. 2b comes from black hole entropy. Gibbons and Hawking Gibbons and Hawking (1977) compute the entropy of the Schwarzschild black hole by computing the action of the Euclidean Schwarzschild metric (which comes entirely from the Gibbons-Hawking-York boundary term York (1972); Gibbons and Hawking (1977)). And, because they compute this action for the one-sided (blue) spacetime, they obtain a non-zero answer. This is a clue that the one-sided Euclidean metric is physically incomplete. If, instead, we compute the action/entropy for the two-sided Euclidean spacetime, the answer vanishes (as the two boundary terms, on isometric but oppositely oriented boundaries, make equal and opposite contributions, see Appendix D). This suggests a picture in which the full (blue+red) geometry corresponds to the full (pure, zero entropy) quantum state, while the half (blue) geometry corresponds to the mixed state obtained by tracing out the other (red) half; and the black hole entropy corresponds to the entanglement entropy (i.e. the von Neumann entropy of the blue side, after tracing out the red side). This in turn explains why the entropy is proportional to the area of the horizon (since the entanglement entropy between two regions is characteristically proportional to the size of the boundary separating them).
If we Wick rotate the full Euclidean metric (3) back to Lorentzian signature using $\tau=it$, we obtain a metric that now covers the full exterior $r>2m$ in Fig. 1a (i.e. regions $I$ and $I^{\prime}$). This is the Einstein-Rosen bridge Einstein and Rosen (1935), which is incomplete as it stands, and needs to be extended to a complete spacetime. The conventional completion is the Kruskal extension of Fig. 1; but here we suggest a different completion. The simplest way to introduce this alternative possibility is to extend the metric across the BH (WH) by replacing $t$ with the ingoing (outgoing) Eddington-Finkelstein time coordinate $v_{+}$ ($v_{-}$), related by $dv_{\pm}=dt\pm\frac{dr}{f(r)}$, so that the metric becomes
@@
Note that the ingoing coordinate $v_{+}$ smoothly extends exterior $I$ onto the upper BH, and exterior $I^{\prime}$ onto the lower BH (Fig. 3a); and for this extension to be well defined (to not assign the same $(v_{+},\sigma)$ coordinate to two distinct points) we must identify antipodal points on the upper and lower BH horizons, as shown in Fig. 3c (e.g. $P\sim P^{\prime},Q\sim Q^{\prime}$). Similarly, the ingoing coordinate $v_{-}$ extends $I$ onto the lower WH and $I^{\prime}$ onto the upper WH (Fig. 3b), and we must identify antipodal points on the two WH horizons (e.g. $R\sim R^{\prime},S\sim S^{\prime}$). [^2] Note that we do not identify all antipodal points in Fig. 1a, as has sometimes been considered in the past Gibbons (1986); in particular, we do not identify points in exterior $I$ with points in exterior $I^{\prime}$. So on the “other side” of the BH horizon from exterior $I$ is exterior $I^{\prime}$ (rather than interior $II$); and similarly for the WH horizon. This has glued exterior $I$ directly to exterior $I^{\prime}$, and removed interiors $II$ and $II^{\prime}$ altogether (Fig. 3c).
Figure 3: (a) $(v_{+},\sigma)$ coordinates; (b) $(v_{-},\sigma)$ coordinates; (c) the Penrose diagram for the eternal black mirror: exteriors I, I’ are glued together, BH to BH and WH to WH, as indicated (e.g. $P\sim P^{\prime}$, $R\sim R^{\prime}$), and with antipodal inversion of the $(\theta,\varphi)$ 2-sphere (see Appendix B), so the interiors II, II’ and their curvature singularities (dark grey regions) are removed.
The resulting spacetime is the simplest (Schwarzschild) black mirror: eternal, static, uncharged. It is described by the metric (5), and covered by two coordinate patches, which together form a complete atlas: the ingoing patch ($-\infty<v_{+}<\infty,-\infty<\sigma<\infty$) which covers everything except the WH horizon, and the outgoing patch ($-\infty<v_{-}<\infty,-\infty<\sigma<\infty$) which covers everything except the BH horizon. In these coordinates, all components of the metric $g_{\mu\nu}$ and Riemann tensor $R^{\alpha}_{\;\;\beta\gamma\delta}$ (and all curvature invariants, such as the Kretschmann scalar) are everywhere smooth, analytic and finite, and the vacuum Einstein equations $R_{\mu\nu}=0$ are satisfied.
Note that the Lorentzian and Euclidean geometries are now in perfect correspondence: in both, the $\sigma=0$ surface (i.e. the origin of the double cigar in Euclidean signature, or the event horizon in Lorentzian signature) serves as the boundary between two mirror-image exterior regions, with no interior regions or curvature singularities.
But there is no free lunch! Although we have gotten rid of the black hole’s curvature singularity, it has been replaced by another (milder) singularity. In particular, if we compute the eigenvalues of $g_{\mu\nu}$ (in $v,\sigma$ coordinates) then, in addition to the angular eigenvalues $\lambda_{\theta}=r^{2}(\sigma)$ and $\lambda_{\varphi}=r^{2}(\sigma){\rm sin}^{2}\theta$, we see the remaining two (positive and negative) eigenvalues
@@
Figure 4: The Penrose diagram for: a black hole formed by gravitational collapse (a); a black mirror formed by gravitational collapse, in unfolded view (b) or folded view (c); a black hole which collapses, then evaporates (d), improved to show shrinking of horizon due to Hawking evaporation (e); a black mirror which collapses, then evaporates (f). Some infalling particle trajectories are also shown, see Appendices B, C.
@@ III Discussion
Figure 5: Asymptotically flat spacetime with CPT-symmetric boundary conditions: states $|\bar{i}\rangle$ and $|\bar{f}\rangle$ are the CPT conjugates of states $|i\rangle$ and $f\rangle$, respectively (see Appendix B).
So far, we have considered the general stationary black mirror. Now let us explain how we think the general time-dependent case works. (Since this part of the story can no longer be described explicitly analytically, it is correspondingly less certain and more speculative.) We start with the Penrose diagram in Fig. 4a for an asymptotically-flat black hole formed by realistic gravitational collapse, with a blue curve showing the trajectory of an infalling particle (the arrow shows the direction of flow of electric charge). The corresponding black mirror is again obtained by identifying two CPT mirror copies of the exterior at their respective horizons, as in Fig. 4b: now we see the infalling blue trajectory’s smooth analytic extension (the red trajectory, see Appendix C). For visualization, it is convenient to fold the diagram in Fig. 4b to obtain the diagram shown in Fig. 4c. In this folded picture, time runs in its usual upward direction on both sheets, so we can explicitly see that both sheets share a common future null boundary (i.e., a BH), and the previous trajectory may now be reinterpreted as the (blue) particle falling into the BH from one side annihilating with its CPT mirror (red) anti-particle falling in from the other. (The arrow flip is explained in Appendix B.)
In the stationary case, the event horizon and apparent horizon coincide. More generally, where they do not, we conjecture the correct matching surface is the apparent horizon (which is well-defined to linear order in perturbations around the stationary solution Hollands et al. (2024)) since: (i) it is defined locally (unlike the event horizon); and (ii) it appears to be the surface associated to dynamical black hole entropy Hollands et al. (2024) which, as argued above, is also associated to the surface separating the two black mirror exteriors.
@@
i) It is natural to consider the path integral for our universe with CPT-symmetric boundary conditions as shown in Fig. 5 (see Appendices B, D). Whereas the black hole (Fig. 1a) and black mirror (Fig. 1b) both satisfy these boundary condition, the black mirror seems to be the relevant saddle point as it (i) is free of curvature singularities, and (ii) has no additional boundaries (whereas the black hole solution has additional boundaries which are singular and also require further data to be specified).
@@
iii) The black hole information paradox Hawking (1976) and firewall paradox Almheiri et al. (2013) come from imagining that quantum information falls into a black hole, and asking how it can possibly re-emerge (since the black hole interior is causally disconnected from future null infinity, see Figs. 4a, 4d). The black mirror, with no interior, suggests a resolution: when a (blue) particle and (red) anti-particle meet at the horizon, they annihilate and the information is then confined to the horizon (classically, Figs. 4b, 4c), or gradually leaks off the horizon (quantumly), until the black hole evaporates (Fig. 4f).
Figure 6: The Penrose diagram for: (a) an unperturbed CPT-symmetric (two-sheeted) universe; (b) a perturbed CPT-symmetric universe in which two black mirrors have formed.
iv) As a black hole settles to stationarity, we expect it to approach a thermal state, which should also be CPT-symmetric Sewell (1982). The black hole (Fig. 4a) does not do this (it settles down to a spacetime with a BH and no WH, with no way for CPT to act), but the black mirror (Fig. 4b, 4e) does (see Appendix B). (Hawking also argued that CPT conflicts with black hole interiors Hawking (2014).)
v) In the standard picture, one can drop global net charge into a black hole, which is then erased when the hole evaporates, leading to the lore that quantum gravity violates global symmetries Hawking (1975); Witten (2018); Harlow and Ooguri (2019, 2021). By contrast, in the black mirror, global charge falling in on one side is canceled by anti-charge falling in on the other; so evaporation does not imply violation of global symmetry.
vi) This CPT-symmetric picture for black mirrors fits naturally in the CPT-symmetric picture of the cosmos advocated in Boyle et al. (2018, 2022a); Boyle and Turok (2021a, b); Boyle et al. (2022b); Turok and Boyle (2024); Boyle and Turok (2024); Turok and Boyle (2023) (and Kumar (2024)) (see Fig. 6): the two mirror-image sheets of spacetime split at the Bang, and merge at BHs. As before, it is natural to regard the two-sheeted spacetime as representing a pure state, while one sheet represents the mixed state obtained by tracing over the other sheet. This neatly explains why the corresponding entropy, associated with the boundary surface separating the two sheets, is given by the horizon area (in the BH case) Bekenstein (1972, 1973); Gibbons and Hawking (1977), and the total comoving spatial volume (in the cosmological case Turok and Boyle (2024); Boyle and Turok (2024)).